[Hallc_running] [New Logentry] Hall C RC Update: Wednesday, July 10

gaskelld at jlab.org gaskelld at jlab.org
Wed Jul 10 10:45:02 EDT 2019

Logentry Text:
Beam came back yesterday on swing shift, and we re-started data taking just before the start of owl shift.  At about 5 am, we completed the SHMS=5.75 degree setting. We then asked MCC to perform a beam energy measurement (result 3660.2 +/- 2.6 MeV, https://logbooks.jlab.org/entry/3701511).  After completion of the energy measurement, we went to Controlled Access to rotate the SHMS to 6.88 degrees,

The near term plan is to continue with the Q2=0.375 middle epsilon data taking with no planned interruptions. Until the pass change in a couple days.

There will be an RC meeting today at 4:30 pm in the 2nd floor Counting House meeting room. Bluejeans link: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3701706
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190710/66eb87fb/attachment.html>

More information about the Hallc_running mailing list